APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday January 13, 2008 (MJD 54479) ---=== OBSERVING TEAM ===--- Dmitry Bizyaev Dan Oravetz ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- We lost the first half of the night to clouds. Finished cartridge 8 (seguefaint170). Finished cartridge 6 (chunk153). Finished cartridge 4 (chunk153). Acquired signal on cartridge 5 (chunk153). ---=== OBSERVING LOG ===--- Afternoon: ---------- No problems to report. Equipment checked and working properly. No adjustments made to spectro collimation. During the fiducial checks, the only altitude fiducial registered when we slewed from 30 degrees to zenith was the 75 degree fiducial. During the slew down, the zenith fiduical (~87 degrees or so) and the 45 degree fiducial were missed. The manipulator arm interlocks logic is still OFF. 23:15Z - Began ventilation. Doors and louvers opened, fans on. Night: ------ 00:30Z - Enclosure off. Mounted cartridge 9. We did not start setup or go to the field because of clouds. We decided to let the telescope equilibrate while we monitored the weather situation. 01:15Z - The clouds did not clear out and we saw signs of approaching radar returns from the west, so we put the enclosure back on, but left everything ventilating. 06:30Z - The clouds mostly cleared out and the radar returns disappeared, so we decided to open. 06:35Z - Enclosure off. Mounted cartridge 8. Cartridge 8, plate 2867, seguefaint170 ************************************** 06:49Z - gotoField. Guide stars found on gotoField. The guider data window did not recognize most of the guide stars to begin with because the seeing was poor. The DIMM was reporting 2.1" at the time. 07:04Z - Pre-Calibs. 07:12Z - 30-minute exposure. During this exposure, guide stars were sometimes not recognized by the guider data window, but we could see the stars in the guide fibers. The efficiency jumped around between ~20% when optically thin clouds moved through the field, and ~50% even when things appeared clear. 07:46Z - 30-minute exposure. We had clouds move through the field at times during this exposure as well. 08:20Z - Post-Calibs. 08:29Z - 30-minute exposure. Seeing: 2.4" - 3.1" fwhm. This plate is DONE. Cartridge 6, plate 2603, chunk153 ********************************* 09:08Z - gotoField. Guide stars found on gotoField. The wind speed picked up and was around 25 mph out of the northeast. The seeing was poor at ~3.2" fwhm. 09:20Z - Pre.Calibs. 09:30Z - 25-minute exposure. 09:58Z - 25-minute exposure. Seeing: 2.3" - 3.2" fwhm. SOS requested post-cal's after the last science exposure, but we had already switched cartridges. This plate is DONE. Cartridge 4, plate 2646, chunk153 ********************************* 10:32Z - gotoField. Guide stars found on gotoField. 10:37Z - Pre-Calibs. 10:44Z - 25-minute exposure. 11:12Z - 25-minute exposure. 11:41Z - 28.3-minute exposure. Seeing: 1.9" - 2.4" fwhm. This plate is DONE. Cartridge 5, plate 2612, chunk153 ********************************* 12:21Z - gotoField. Guide stars found on gotoField. 12:28Z - Pre-Calibs. 12:35Z - 20-minute exposure. Seeing: 2.5" - 3.2" fwhm. 13:05Z - Enclosure on, doors and louvers closed, fans off. 13:12Z - endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7281 22:11Z 22:35Z 100 O 154.70 160.80 55 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 48910, plate -9999 ------- 22:13 48910 0.0 bias 22:28 48911 0.0 bias ----- sequence 48912, plate 2603 ------- 22:43 48912 10.0 flat 22:45 48913 2.0 arc 22:52 48914 3.0 arc 22:55 48915 3.0 arc ----- sequence 48916, plate 2867 ------- 92.1 50.5 96.5 72.4 07:06 48916 10.0 flat 07:09 48917 2.0 arc 07:44 48918 1800.0 target 3.4 2.7 3.6 4.0 08:19 48919 1803.0 target 2.9 2.0 2.9 2.7 08:22 48920 10.0 flat 08:24 48921 2.0 arc 09:02 48922 1800.0 target 5.0 2.9 5.6 4.8 ----- sequence 48923, plate 2603 ------- 27.1 17.0 31.4 25.8 09:23 48923 10.0 flat 09:25 48924 2.0 arc 09:57 48925 1501.0 target 4.0 2.1 4.3 3.6 10:25 48926 1501.0 target 5.6 3.2 6.3 5.2 ----- sequence 48927, plate 2646 ------- 28.5 15.2 28.5 16.5 10:40 48927 10.0 flat 10:42 48928 2.0 arc 11:12 48929 1501.0 target 9.3 4.5 8.8 5.7 11:40 48930 1501.0 target 9.4 5.2 9.5 5.8 12:11 48931 1702.0 target 9.8 5.4 10.2 6.0 ----- sequence 48932, plate 2612 ------- 3.3 3.1 2.8 2.1 12:30 48932 10.0 flat 12:33 48933 2.0 arc 12:58 48934 1200.0 target 3.3 3.1 2.8 2.1 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 07:23Z 8 2867 49.06 0.0036 45.43 -0.0006 219.43 0.0286 0.999850 09:33Z 6 2603 80.59 0.0048 45.95 0.0002 240.15 0.0228 1.000000 10:46Z 4 2646 60.50 0.0063 71.50 -0.0003 232.36 0.0402 1.000000 12:36Z 5 2612 -34.26 0.0079 74.76 -0.0015 149.75 0.0246 1.000020 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 07:23Z 8 2867 0.99985 1481 2672 250 49 45.4 -7.1 18 77 ---- 3.1" 09:33Z 6 2603 1.00000 0 1465 270 81 45.9 -7.8 23 71 ---- 2.9" 10:45Z 4 2646 1.00000 0 1399 270 61 71.4 -7.8 20 70 ---- 1.9" 12:36Z 5 2612 1.00002 -198 1231 270 -34 74.8 -7.8 20 70 ---- 2.5" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:21Z 25 3 1 215 (SW) 304 - mostly cloudy 22:52Z 29 -3 3 280 (W) 135 - partly cloudy 23:24Z 27 -3 2 257 (WSW) 183 - partly cloudy 23:56Z 27 -1 4 255 (WSW) 168 - partly cloudy 00:28Z 26 1 0 94 (E) 266 - mostly cloudy 01:01Z 25 1 3 90 (E) 249 - mostly cloudy 01:33Z 25 1 4 108 (ESE) 233 - clouds 02:05Z 23 6 7 90 (E) 294 - clouds 02:37Z 24 5 6 163 (SSE) 264 - clouds 03:09Z 24 5 8 124 (SE) 287 - clouds 03:42Z 23 7 10 84 (E) 273 - clouds 04:14Z 21 7 13 89 (E) 326 - clouds 04:46Z 21 7 19 85 (E) 362 - clouds 05:18Z 20 6 21 76 (ENE) 322 - clouds 05:50Z 19 6 26 79 (E) 324 - clouds 06:22Z 19 5 18 85 (E) 299 - partly cloudy 06:53Z 18 6 23 75 (ENE) 260 - partly cloudy 07:26Z 19 7 16 80 (E) 379 - partly cloudy 07:56Z 19 6 22 74 (ENE) 316 - partly cloudy 08:26Z 18 6 21 73 (ENE) 325 - clear 08:57Z 18 5 18 58 (ENE) 259 - clear 09:27Z 17 4 23 63 (ENE) 225 - clear 09:58Z 17 5 23 72 (ENE) 312 - clear 10:29Z 18 5 28 72 (ENE) 344 - clear 11:00Z 18 3 26 72 (ENE) 328 - clear 11:31Z 18 1 23 65 (ENE) 198 - clear 12:01Z 18 6 23 73 (ENE) 307 - clear 12:32Z 18 7 20 77 (ENE) 408 - clear 13:02Z 18 8 20 75 (ENE) 422 - clear ---=== TELESCOPE STATUS ===--- 23:15Z - Began ventilation. Doors and louvers opened, fans on. 00:30Z - Enclosure off. 01:15Z - Enclosure on, but still ventilating. 06:35Z - Enclosure off. 13:05Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 13:11Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 5 Counterweights at: 270 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 83.60 lbs, 23.45 psi Imager 175.10 lbs, 19.21 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_37 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2170 -5.2160 Axial B -0.2142 -0.2142 Axial C -0.0016 -0.0016 Trans D 11.8530 11.8540 Lateral E 10.0711 10.0584 Lateral F 11.6586 11.6586 GALILS Commanded: 4600. -2000. 2650. -12450. -6500. -6550. Actual: 4600. -2000. 2650. -12439. -6477. -6546. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 100.00 -200.00 PriOrient: 0.00 0.00 0.00 99.59 -200.69 SECONDARY: ---------- Focus: 0.00 Air Temp.: -2.7 Alt.: 30.003178 MIGS TONIGHT NOMINAL Axial A 0.8024 0.8025 Axial B 0.6122 0.6120 Axial C 0.5475 0.5474 Trans D 0.6005 0.6008 GALILS Commanded: 1627466. 1653129. 1470862. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.97 60.00 -4.00 -19.92 119.29 ---=== PROBLEMS IN DETAIL ===--- Spectro collimation with cartridge 6 ------------------------------------ 'mechSend s' before: sp1> -17406 -17386 -18492; 2.4C sp2> -9085 -10040 -10184; 0.3C IDL 5.5a> apofix, 48913, /test sdR-b1-00048913 QUALITY=[excellent] -> ['test'] sdR-b1-00048913 QUALITY=[excellent] -> ['test'] sdR-b2-00048913 QUALITY=[excellent] -> ['test'] sdR-b2-00048913 QUALITY=[excellent] -> ['test'] sdR-r1-00048913 QUALITY=[excellent] -> ['test'] sdR-r1-00048913 QUALITY=[excellent] -> ['test'] sdR-r2-00048913 QUALITY=[excellent] -> ['test'] sdR-r2-00048913 QUALITY=[excellent] -> ['test'] File sdHdrFix-54479.par contains 2 declared changes. APOFIX: Trying to lock the logfile /data/spectro/spectrologs/54479/logfile-5447 9.fits APOFIX: Reading the logfile /data/spectro/spectrologs/54479/logfile-54479.fits MRDFITS: Binary table. 25 columns by 4 rows. % READFITS: FITS header has NAXIS or NAXISi = 0, no data array read MRDFITS: Binary table. 19 columns by 8 rows. MRDFITS: Binary table. 24 columns by 4 rows. MRDFITS: Binary table. 25 columns by 4 rows. MRDFITS: Null image, NAXIS=0 MRDFITS: Binary table. 6 columns by 16 rows. MRDFITS: Extension past EOF IDL 5.5a> collimate, 48914 Reading /data/spectro/54479/sdR-b1-00048914.fit Reading /data/spectro/54479/sdR-b1-00048915.fit COLLIMATE: Min offset = -0.73 pix COLLIMATE: Max offset = 1.51 pix COLLIMATE: RMS across CCD = 0.61 pix COLLIMATE: Mean offset = 0.40 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -14.3 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-54479-b1-00048914.log PostScript file = Collimate-54479-b1-00048914.ps Reading /data/spectro/54479/sdR-b2-00048914.fit Reading /data/spectro/54479/sdR-b2-00048915.fit COLLIMATE: Min offset = -0.15 pix COLLIMATE: Max offset = 0.63 pix COLLIMATE: RMS across CCD = 0.18 pix COLLIMATE: Mean offset = 0.29 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -10.5 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-54479-b2-00048914.log PostScript file = Collimate-54479-b2-00048914.ps Files not found Files not found % Program caused arithmetic error: Floating illegal operand No adjustments made to spectro collimation.